### v1

# AutoV: Learning to Retrieve Visual Prompt for Large Vision-Language Models

Abstract
Inspired by text prompts in large language models (LLMs), visual prompts have been explored to enhance the reasoning capabilities of large vision-language models (LVLMs). Current methods design heuristic visual prompts, such as overlaying a text-query-guided attention heatmap on the original input image. However, designing effective prompts manually is challenging and time-consuming, and it often fails to explore the benefits of different visual prompts, leading to sub-optimal performance. To this end, we propose AutoV that learns to automatically select the optimal visual prompt from various candidates based on given textual queries and the input image. To train AutoV, we developed an automatic data collection and labeling pipeline that evaluates various visual prompts with a pre-trained LVLM. We input a set of visual prompts into the LVLM and rank them according to the prediction losses generated by the model. Using the ranking as a supervision signal, we train AutoV to automatically choose the optimal visual prompt from various visual prompts for LVLMs. Experimental results indicate that AutoV enhances the performance of various LVLMs across multiple popular image understanding tasks. For instance, LLaVA-OV with AutoV achieves accuracy gain on LLaVA, and AutoV boosts Qwen2.5-VL by on MMMU, highlighting its potential as an optimal visual prompting method for LVLMs.
Shanghang Zhang at
1 Introduction
Recent advancements in large language models (LLMs) [39, 10, 1, 47, 37, 8] have spurred significant progress in large vision-language models (LVLMs), which connect visual encoders with language model decoders to effectively adapt textual reasoning capabilities to visual understanding tasks [31, 32, 4, 13, 26, 25, 5, 24]. Therefore, visual input plays a crucial role in LVLMs, enabling precise perception and fine-grained grounding across various multimodal scenarios.
To further improve the perception and reasoning of LVLMs, visual prompting [55, 30, 53, 43] has emerged as an important technique analogous to textual prompting in LLMs [22, 65, 49]. Visual prompts enhance model performance by adding visual annotations such as masks, colorful circles, or heatmaps, thus guiding the attention of LVLMs toward crucial image regions of interest. However, existing visual prompting methods predominantly rely on heuristically designed prompts, limiting their ability to fully exploit the inherent diversity of visual prompts. As illustrated in Figure 1, we list the benchmark-level performance of existing heuristic prompt methods (e.g., FGVP [53], Red Circle [43], and API [55]). While all three methods demonstrate effectiveness on logic-intensive MMMU tasks, both blur-based FGVP and bounding prompts adversely affect vision-centric benchmarks (e.g., VizWiz [9]) and OCR-dependent tasks (e.g., TextVQA [44]). This is because heuristic prompt methods fail to automatically tune hyperparameters for each benchmark or even each instance.
Hence, to develop a more effective visual prompt strategy suitable for diverse vision-language tasks, it is vital to adaptively retrieve the optimal visual prompt that most significantly improves the perception and reasoning of LVLMs in response to specific queries. Motivated by this, we introduce AutoV, an innovative visual prompting retrieval framework designed to dynamically select the optimal visual prompt from a diverse set of candidates, explicitly tailored to the textual query and input image. Specifically, we first generate a diverse set of visual prompt candidates111In this work, we use attention-based visual prompts from various transformer layers of CLIP in API [55]. and encode their representation. This diversity enables the framework to capture distinct visual representations that may be optimal for various query-image pairs. Next, we design a lightweight ranking network that efficiently integrates visual and textual information to predict the rankings preferences over prompt candidates. Finally, the highest-ranking candidate is chosen as the input for LVLMs during inference.
To train the ranking network, we establish an automated pipeline for collecting and labeling training data, evaluating the rank of each visual prompt candidate based on the prediction losses of a pre-trained LVLM. Subsequently, the ranking network is supervised using a reward loss that differentiates between selected and rejected candidates. This strategy offers reliable supervision, allowing AutoV to attain adaptive, query-specific prompt retrieval without the need for extensive manual annotation.
AutoV achieves significant performance improvements across various prominent LVLM architectures and diverse image understanding benchmarks, including fine-grained comprehension, real-world visual assistance, and complex open-ended reasoning tasks. For instance, integrating AutoV with LLaVA-OneVision and InternVL2 leads to improvements of and on LLaVA, respectively. Additionally, AutoV enhances Qwen2.5-VL on MMMU by 1.9%, highlighting its significant advantage in adaptive visual prompting retrieval. In summary, our contributions are as follows:
We present AutoV, a novel automatic visual prompting framework for large vision-language models that adaptively retrieves the optimal visual prompt from a diverse candidate set, explicitly tailored to textual query and visual context pairs.
We create a scalable and automated data collection pipeline equipped with a reward-based loss to effectively train a lightweight ranking network, facilitating accurate and adaptive visual prompt selection without the need for extensive manual annotations.
We perform comprehensive experiments to validate the efficacy and generalizability of AutoV across various multimodal models and benchmarks, showcasing consistent and significant performance improvements over existing visual prompting methods. Once trained, AutoV integrates seamlessly into various LVLMs without additional fine-tuning.
2 Related Work
Large Vision-Language Models. Recent advances in large language models [39, 10, 1, 47, 37, 8] have catalyzed a new wave of large vision-language models [31, 2, 4, 13, 26, 25]. By coupling a high-capacity visual encoder with an LLM decoder, LVLMs transfer textual reasoning to visual understanding, and the quality of input images determines the lower bound of multimodal abilities. For instance, the popular LLaVA-OneVision [24] and Qwen2.5-VL [5] adopt a higher and dynamic resolution recipe for training to enhance their fine-grained reasoning, while vision-centric scaling in InternVL2 [11] enlarges the ViT [15] backbone to boost grounding capability. However, the above methods require training times of several hundred GPU days. Here, can we further boost the existing LVLMs by optimizing the visual input itself during the inference stage? The answer is Yes. Visual prompting [45, 29] is an emerging and effective technique in this direction.
Visual Prompting for LVLMs. LVLMs relying solely on textual inputs exhibited two key limitations: visual hallucinations [6, 18] and language-driven biases [50, 38]. To better mitigate these challenges, visual prompting [60, 45, 29, 51] has emerged as an alternative way of supplying cues in the visual modality. Compared with textual prompts, it is more concise and explicit, while allowing guidance that reaches fine-grained, pixel-level details. Early work [43] shows that simply drawing a red circle around a target object steers CLIP [40] attention, yielding great performance on referring-expression and keypoint-localization benchmarks. FGVP [53] extends this idea with fine-grained prompting, replacing coarse shapes with blur-reversal prompts based on segmentation masks, further improving accuracy on RefCOCO [61]. AlphaCLIP [45] trains an augmented version of CLIP with an auxiliary alpha channel for indicating regions at the expense of increased training overhead. API [55] proposes attention prompting on images, overlaying text-guided attention heatmaps, boosting LVLMs on VQA, and open-ended reasoning tasks. Although previous studies have explored various visual prompts, it remains unclear which prompt is optimal for a specific image-question pair, leading to a decrease in the practicability of visual prompts. To address this limitation, we introduce our visual prompting method that learns to retrieve the optimal visual prompt for a specific image-question pair.
Visual Retrieval for LVLMs. A variety of strategies have been proposed for visual retrievals, such as visual token selection [12, 63, 54] for LVLMs. Existing visual prompt selection pipelines generally adhere to three main paradigms. (1) The heuristic paradigm identifies benchmark-level optimal parameters through extensive evaluation but lacks instance-specific adaptation. (2) The random paradigm offers a stochastic baseline characterized by instance-level variation, but it lacks performance guarantees. This establishes it as the lower bound for adaptive approaches. (3) The Mixture-of-Experts (MoE) paradigm [16, 7, 66] provides the most direct form of adaptation by utilizing the GateNet specifically for visual retrieval. However, MoE gates are susceptible to overfitting during training, which can result in poor generalization. To mitigate these limitations, we propose an efficient ranking network specifically designed for visual retrieval, which models the pairwise preference ordering.
3 Proposed Approach: AutoV
In this section, we present AutoV for optimal visual prompt retrieval. Our framework includes four critical components: (1) feature extraction of prompt candidates, (2) candidate ranking consisting of modality interaction and single-modality mapping modules, (3) reward loss supervision along with automated data generation, and (4) an efficient inference pipeline. The overall architecture is illustrated in Figure 2, where the four modules are color-coded for clarity. Best viewed in color.
3.1 Feature Extraction of Prompt Candidates
The primary goal is to effectively choose the most appropriate visual prompt in response to various text queries and input images. For a group of input visual prompts , we employ a visual encoder (e.g., like CLIP [40]) to generate the visual feature:
Next, we apply a projection matrix to transform each visual feature into language embedding tokens , aligning the same dimensionality as the word embedding space in the language model. As a result, we obtain a set of visual candidates represented in the form of visual tokens to rank, where denotes the length of vision tokens and is the language embedding dimension. Notably, both the visual encoder and the alignment projector are directly inherited from a fully trained LVLM, and we adopt them without any additional fine-tuning.
3.2 Candidate Ranking Network
The goal of the ranking network is to effectively rank the visual prompt candidates based on their relevance to the textual query, by modeling the relationship between the visual tokens and the input text. Therefore, we first propose the modality interaction module to enable visual tokens to fuse the context information of text tokens. Inspired by observations in [62] that modality-fusion occurs in the early layers of the LLM, we use the first layer of the LLM decoder for modality interaction. Thus visual candidate tokens and text tokens are concatenated and fed into the interaction module:
with the outputs are and . Next, we input visual candidate tokens into a vision mapping module, which is implemented using a feed-forward network (FFN). The generated final candidate features are , where is the output dimension of the vision mapping module to save computational cost. Besides, the query tokens are also converted by a text mapping module to get the query embeddings . In AutoV, we define visual retrieval as a customized ranking task rather than a classification or regression problem, and the reason is discussed in Section 3.3. Therefore, we calculate the reward scalar for given visual candidate tokens , where the context similarity with the query serves as the quantitative target. We utilize a cross-attention operation to achieve:
Here, only the vision and text mapping modules need additional training, each consisting of a single FFN with two linear layers. The total parameters are , resulting in lower training costs.
3.3 Reward Loss Supervision
Automated Data Generation. We create customized data to train the ranking network to select the optimal visual prompt for a given LVLM. Our approach is grounded in a simple intuition: a superior visual prompt should yield lower language modeling loss when input into any trained LVLM. We optimize the ranking network using relative loss instead of a one-hot selection vector, as correctness can stem from the LVLM’s language priors without relying on any input content [52, 56]. In contrast, the inference loss indicates the confidence of models in generating answers based on the input visual-question pair: a lower loss indicates higher confidence [17, 64].
For each group of visual candidates, we generate (VP, query) pairs and comprehensively evaluate them. Each pair is processed by a fully trained vision-language model to compute its combination loss, defined as the modeling loss relative to the ground truth, and sorted to get a rank. To ensure high-quality ranking data, we implement two filtering criteria for (VP, query) pairs: (1) we remove pairs exhibiting low loss variance, as their insensitivity to visual prompts suggests answers are generated from the language priors rather than actual input content; (2) we exclude pairs with excessively high average loss, which indicates outlier case or weak visual prompt-query correlation.
For each raw (image, query) pair, we generate a training sample structured as a quadruple:
with an illustrative example provided in Figure 2, and the dataset we adopt is detailed in Appendix B.
Reward Loss Supervision. Inspired by the reward modeling from human feedback in OpenAI [36], we train our ranking network using reward modeling to align its partial order with the optimal visual prompt. Here, we model visual retrieval as a customized ranking task rather than a classification or regression problem. This is because the objective is not to assign absolute labels or scores, but to identify the most appropriate prompt from a set of candidates by comparing their relative effectiveness in enhancing multimodal understanding. Ranking allows the model to learn fine-grained preferences across diverse query-image combinations and better align with the real-world setting.
In Figure 3, we pair all visual prompt candidates exhaustively, ensuring each prompt is combined with every other prompt exactly once. This results in generating training pairs for each instance. The reason is that, unlike pointwise scoring, which assigns low scores to all responses for inherently challenging tasks, pairwise comparison allows for finer discrimination between responses. Based on the rank and loss from the quadruples, we reinforce the visual prompt with lower loss (chosen sample) for each pair, while penalizing the higher-loss visual prompt (rejected sample).
Specifically, the reward loss is designed as:
where is the scalar output of the ranking network for the chosen visual prompt out of the pair, represents the rejected one, and denotes the Sigmoid function.
3.4 Efficient Inference Pipeline
During the inference stage, we input several suitable visual prompt candidates along with the query into the large vision-language model and then pass them through our ranking network, as illustrated in Figure 2. Next, the ranking network outputs the corresponding reward scalar of each visual prompt. We then select the highest-scoring visual prompt, concatenate its visual tokens with text embeddings, and decode their representation through the LLM to produce the final answer.
4 Experiments
4.1 Experiment Settings
Models. We apply our AutoV to four mainstream open-source large vision-language models, i.e., LLaVA-v1.5 [33], LLaVA-OneVision [24], Qwen2.5-VL [5], and InternVL2 [14]. In particular, LLaVA-v1.5 employs CLIP-pretrained ViT [40] as the visual tower and Llama2 [48] as the LLM backbone. LLaVA-OneVision enhances this design with SigLIP [59] for high-resolution refinement and replaces the LLM with Qwen2 [3]. Qwen2.5-VL features a dynamic resolution encoder, while InternVL2 adopts InternViT [13] for visual encoding and InternLM2 [11] as its LLM decoder.
Benchmarks. We evaluate AutoV on six widely-adopted image understanding benchmarks, including MMMU [58], POPE [27], VizWiz [9], MMVet [57], VQA (TextVQA) [44], and LLaVA (LLaVA-Bench in-the-wild) [31]. The evaluation benchmarks contain various aspects: hallucination suppression (POPE), STEM reasoning (MMMU), real-world visual assistance (VizWiz), fine-grained comprehension (MMVet), text-rich understanding (VQA), and open-world reasoning (LLaVA).
4.2 Main Results
| Method | MMMU | POPE | VizWiz | MMVet | VQA | LLaVA | Avg. |
| LLaVA-v1.5 (7B) | |||||||
| Base Model | 36.3 | 85.9 | 50.0 | 30.6 | 58.2 | 65.4 | |
| FGVP (NIPS23) | 36.9 | 72.6 | 47.5 | 29.0 | 46.3 | 55.6 | |
| Circle (ICCV23) | 37.0 | 84.5 | 49.6 | 31.3 | 55.5 | 62.2 | |
| API (ECCV24) | 37.4 | 86.0 | 50.3 | 31.8 | 57.7 | 65.8 | |
| AutoV | 38.3 | 86.8 | 51.0 | 32.8 | 58.8 | 66.6 | |
| LLaVA-v1.5 (13B) | |||||||
| Base Model | 36.7 | 85.9 | 53.6 | 36.1 | 61.2 | 66.6 | |
| FGVP (NIPS23) | 36.9 | 66.8 | 46.8 | 29.7 | 49.2 | 63.7 | |
| Circle (ICCV23) | 36.9 | 84.6 | 53.8 | 31.2 | 58.1 | 66.9 | |
| API (ECCV24) | 37.3 | 85.6 | 54.6 | 35.6 | 60.6 | 67.6 | |
| AutoV | 38.5 | 86.9 | 55.4 | 37.0 | 61.9 | 68.5 | |
| LLaVA-OneVision (7B) | |||||||
| Base Model | 49.4 | 89.1 | 58.2 | 48.8 | 76.1 | 80.6 | |
| FGVP (NIPS23) | 48.6 | 82.9 | 52.0 | 45.5 | 65.4 | 70.8 | |
| Circle (ICCV23) | 48.4 | 88.2 | 59.6 | 47.2 | 73.0 | 76.9 | |
| API (ECCV24) | 49.0 | 88.9 | 58.7 | 48.1 | 75.9 | 80.9 | |
| AutoV | 50.0 | 90.0 | 59.4 | 49.9 | 77.3 | 82.3 | |
| InternVL2 (8B) | |||||||
| Base Model | 48.6 | 87.8 | 58.6 | 46.5 | 77.0 | 74.6 | |
| FGVP (NIPS23) | 47.2 | 84.2 | 52.3 | 36.8 | 70.7 | 70.2 | |
| Circle (ICCV23) | 48.3 | 86.0 | 57.5 | 40.3 | 74.9 | 74.0 | |
| API (ECCV24) | 48.3 | 87.9 | 58.0 | 46.3 | 77.0 | 75.6 | |
| AutoV | 49.2 | 88.5 | 59.1 | 47.2 | 77.9 | 76.8 | |
| Qwen2.5-VL (7B) | |||||||
| Base Model | 50.1 | 87.5 | 69.5 | 56.6 | 83.0 | 98.0 | |
| FGVP (NIPS23) | 49.6 | 81.6 | 65.8 | 43.1 | 78.0 | 90.4 | |
| Circle (ICCV23) | 48.2 | 84.2 | 66.7 | 44.5 | 80.5 | 89.5 | |
| API (ECCV24) | 51.0 | 87.6 | 69.3 | 56.8 | 82.3 | 98.1 | |
| AutoV | 52.0 | 88.1 | 70.0 | 57.6 | 83.4 | 98.8 | |
Our results are listed in Table 1, and we compare AutoV with three previous visual prompting methods designed for LVLMs, including FGVP [53], RedCircle [43], and API [55]. Performance improvements and decrements listed in the table are calculated relative to the base model. The training data and implementation details can be found in Appendix B and D.
The main observations from the experimental results are as follows: (1) AutoV significantly outperforms existing methods, demonstrating consistent advantages across all models and benchmarks. Specifically, for LLaVA-7B, LLaVA-13B, LLaVA-OneVision, InternVL2, and Qwen2.5-VL, the average improvements relative to the base model are , , , and , respectively. Our method performs particularly well for LLaVA-7B on MMVet, with an improvement of . Excluding it, our method appears more effective and robust on the VQA task, where other methods all fail to achieve positive gain. (2) Our approach shows the most significant performance advantage on MMVet compared to the API. With an average score of for our method versus for theirs, the relative difference of underscores the importance of accurate case-level visual cues retrieval instead of benchmark-level visual prompts selection. The other two visual prompting methods, lacking the ability to adapt to different queries, do not perform well on VQA tasks. (3) Our method further supplements inherently less accurate models. For instance, LLaVA-7B equipped with our approach, achieves higher accuracy on the MMMU and POPE tasks than the 13B-scale model. This means that our method can be utilized for model efficiency on certain specific tasks: a few additional MLP layers are a better choice compared with hundreds of times larger in parameter size. (4) Both Qwen2.5-VL and InternVL2 integrate the retrieval-augmented prompting strategy from AutoV, which is pre-trained on LLaVA-OneVision. This integration results in an average accuracy gain of . Notably, the improvements are consistent and observed in all individual tasks, including complex reasoning (e.g., MMVet) and real-world scenarios (e.g., VizWiz). This demonstrates that the effectiveness of the retrieval-enhanced prompting strategy is model-agnostic and generalizes well across heterogeneous LVLM architectures, regardless of pretraining scales or backbone variations.
4.3 Ablation Study
4.3.1 How does ranking compare to other retrieval methods?
| Method | MMMU | VizWiz | MMVet | Avg. |
| Baseline | 36.3 | 50.0 | 30.6 | 39.0 |
| only | 36.5 | 50.0 | 31.8 | 39.4 |
| only | 36.9 | 50.3 | 31.2 | 39.5 |
| only | 37.4 | 49.9 | 31.7 | 39.7 |
| only | 37.1 | 50.3 | 30.8 | 39.7 |
| Random | 37.2 | 50.4 | 31.6 | 39.7 |
| Mixture-of-Experts | 37.6 | 50.4 | 31.7 | 39.9 |
| AutoV | 38.3 | 51.0 | 32.8 | 40.7 |
To better analyze why ranking is more suitable for retrieving visual prompts, we list three other retrieval methods and provide experimental evidence for comparison, including heuristic retrieval, random retrieval, and mixture-of-experts retrieval. All experiments are conducted on LLaVA-v1.5.
Heuristic Fixed Retrieval. Heuristic retrieval only involves a single fixed type of visual prompts, e.g., red circle [43], or attention image [55], based on its predefined rules or heuristics. Furthermore, to identify the best rules, they need to conduct benchmark-level multiple rounds of trials, consuming a significant amount of reasoning resources. As shown in rows 3 to 6 of Table 2, we find that each visual prompt could potentially be the optimal solution for different benchmarks. For instance, achieves the best results on MMMU, while is optimal on MMVet, and there is a significant variation in the performance of different types of visual prompts across various benchmarks. In summary, heuristic retrieval offers simplicity in retrieval processes with the need for conducting benchmark-level multiple rounds of trials.
Random Retrieval. Random retrieval picks up one visual prompt from the candidate pool without any specific criteria or order. The th row of Table 2 lists its results, comparing with the individual visual prompt, we observe that the random retrieval’s performance is on par with or slightly superior across the three benchmarks. For instance, the random algorithm outperforms in MMMU by , in MMVet by , and even surpasses the best results of the VP candidates in VizWiz. The above experiment demonstrates that by flexibly selecting from multiple visual prompt candidates, even random retrieval has the potential to outperform individual visual prompts, and the conclusion provides valuable insights for our approach.
MoE (or GateNet) Retrieval. Mixture-of-experts (MoE) retrieval is a prevalent technique used to identify the most relevant expert for a model decision-making process, where the Gumbel Softmax function is employed to assign weights to the GateNet in a differentiable manner during training [16]. In our scenario, the features of multiple visual prompts are input to GateNet to determine which one should be retained and fed into the LLM. We insert GateNet into the output of the projection layer and train it with the same amount of data and training time as AutoV. The results can be seen in the th row of Table 2. We observe a improvement with MoE compared to the baseline. However, the improvement compared to the random algorithm was only , lagging significantly behind AutoV by . We suspect that the Mixture-of-Experts architecture encounters a lack of generalization ability during the fine-tuning stage, which can easily lead to overfitting. Additionally, this architecture does not incorporate textual modality for guidance. Therefore, MoE retrieval may not be the optimal approach for visual prompt retrieval in LVLMs. It is important to adopt a more robust framework that integrates textual guidance to enhance overall performance effectively.
Retrieval Visualization. We visualize the prompt distribution of retrieval strategies in practical tasks. Taking the MMVet task as an example, Figure 4 presents a bar graph comparing the retrieval decisions of GateNet and our AutoV. The bar heights indicate the frequency of visual prompt retrieval, while the overlaid line plots depict the independent accuracy for each prompt. Our analysis reveals a notable discrepancy between the GateNet retrieval strategy and the actual inference accuracy of visual prompts. For instance, GateNet retrieves , the second-most useful prompt, only times (28/218, of cases), despite its high accuracy. In contrast, AutoV exhibits stronger alignment with the standalone accuracy of visual prompts, and prefers to select and , the two prompts that own superior accuracy.
4.3.2 More candidates perform better?
| Number | Source | MMMU | VizWiz | MMVet | Avg. |
| 0 | Vanilla image | 36.3 | 50.0 | 30.6 | 39.0 |
| 2 | + {L, L} | 37.3 | 50.3 | 32.0 | 39.9 |
| 3 | + {L} | 37.8 | 50.8 | 32.3 | 40.3 |
| 4 | + {L} | 38.3 | 51.0 | 32.8 | 40.7 |
| 5 | + {L} | 38.2 | 50.8 | 32.8 | 40.6 |
| 6 | + {L} | 38.5 | 50.7 | 32.6 | 40.6 |
| 7 | + {RedCircle} | 38.7 | 50.7 | 32.9 | 40.8 |
| 8 | + {FGVP} | 38.7 | 50.4 | 32.4 | 40.5 |
In AutoV, having more visual prompt candidates can offer a more diverse selection, enhancing the flexibility and adaptability of LVLMs. However, is it always better to have more visual prompts? In the main experiment, we used the attention prompting images [55] from layers , , , and in the CLIP [40] model as inputs, with a total of visual prompts. Here, we consider extra scenarios ranging from prompts and conduct experiments on the MMMU, VizWiz, and MMVet benchmarks. The results are listed in Table 3. First, we observe that when the visual prompts themselves can lead to performance improvement, the accuracy increases as their quantity grows, but the rate of accuracy improvement slows down as the number continues to rise. Taking the experiments on MMMU as an example, as the number of prompts increases, the performance gain remains at when increasing from to prompts. However, when further increasing to prompts, the improvement drops to , even though FGVP alone can contribute a gain. Then, even if a single visual prompt itself degrades accuracy, incorporating it into the candidate pool only marginally harms the overall performance, thanks to the inherent retrieval capability of AutoV. Taking the results on VizWiz as an example: although both RedCircle and FGVP individually show negative gains ( and respectively), including them results in either no overall accuracy drop or just a minimal decrease. In summary, both the quality and quantity of visual prompts impact overall performance. We recommend selecting known high-quality visual prompts as candidates, while limiting to fewer than instances to strike the balance between accuracy and efficiency.
4.3.3 The transferability of AutoV
| Models | VizWiz | LLaVA |
| Gemini-1.5-Pro | 45.9 | 77.9 |
| +AutoV | 55.5 | 81.1 |
| GPT-4o | 52.8 | 68.0 |
| +AutoV | 61.8 | 70.9 |
| Qwen-VL-Max | 64.6 | 85.8 |
| +AutoV | 71.9 | 88.4 |
The retrieval strategy in our AutoV, trained on one LVLM, can be transferred to various other LVLMs. First, our method demonstrates broad applicability across open-source models. In Table 1, both InternVL2 and Qwen2.5-VL achieve a improvement by directly adopting the retrieval strategy from AutoV pre-trained on LLaVA-OneVision. More remarkably, Table 4 reveals that even closed-source models like Gemini-Pro [46] and GPT-4o [20] show substantial gains, averaging and points respectively. Therefore, AutoV owns good transferability, as its retrieval strategy consistently boosts performance across both open- and closed-source LVLMs without model-specific adaptation.
4.3.4 The Efficiency of AutoV
While we enhance LVLM performance through AutoV, we uphold efficiency optimizations to ensure practicality. Initially, training a high-quality AutoV on the LLaVA-7B model within epochs only requires hours utilizing NVIDIA A-GB GPUs. Furthermore, because of its transferability, training additional AutoV instances for each vision-language model is unnecessary, and we can generate the training samples via smaller models. Lastly, although AutoV introduces minimal additional FLOPs during inference (e.g., multiple image pre-processing and visual feature extraction), we can reduce latency by downsizing the final similarity matrix and single-prompt selection before LLM processing. As a result, the inference latency of the LLaVA-7B model increases by just 6.4 ms (from 254.8 ms to 261.2 ms), while yielding a 1.3% gain in average accuracy.
4.3.5 Qualitative Visualization
To further validate the effectiveness of our retrieval strategy, we visualize an example sourced from LLaVA-Bench. As shown in Figure 5, we list the output results of AutoV when we enter different queries. When faced with a question that focuses on the global context information, AutoV chooses where the black mask removes the background. While referring to the "intended effect" in the image, AutoV selects , which directly uses red circles to highlight the weird spots. More visualization retrieval cases can be found in the Appendix G.2.
5 Conclusion
In this paper, we introduced AutoV, an adaptive visual prompting framework designed to effectively retrieve optimal visual prompts tailored to textual queries and image contexts for large vision-language models (LVLMs). By leveraging a lightweight ranking network trained via a reward-based supervision strategy, AutoV significantly surpasses existing heuristic prompting methods. Extensive experiments across multiple LVLM architectures and benchmarks demonstrate consistent improvements. Future research could explore integrating AutoV with more advanced prompt generation techniques or extending its framework to broader multimodal contexts beyond vision-language scenarios.
Appendix
Appendix A Attention Visual Prompts used in Experiments
Alignment in Vision Foundation Models. As vision foundation models advance rapidly [40, 35, 59, 21], pretrained on language-image pairs, they serve as the visual encoders for large vision-language models. A representative model is CLIP [40], , which consists of a vision encoder and a text encoder, mapping the image and text prompt into the aligned latent space. Therefore, we can directly obtain the similarity between vision modality and text modality feature with .
Attention Visual Prompts. Inspired by the idea that text-image attention (or similarity ) can be visualized as visual prompts for LVLMs [55, 30], where similarity heatmaps overlaid on the image highlight key regions requiring greater attention for prompt-based reasoning. In this work, we adopt a fine-grained similarity visualization manner in API [55], with the -th layer similarity defined as follows:
where denotes the number of transformer layers, is the projection layer with a fully-connected layer and normalization operation, and MSA stands for multi-head self-attention structure. is input vision tokens for the -th layer, while indicates the class token in the token sequence .
Appendix B Training Data
Our training source dataset is composed of several components used in LLaVA [31] to ensure that AutoV is trained on a diverse range of image-question pairs. It includes K samples from the COCO train dataset [28], K samples from the Visual Genome (VG) dataset [23], K samples from the GQA dataset [19], K samples from the OCRVQA dataset [34], and K samples from the TextVQA dataset [44]. In total, the source data contains K image-question pairs. Here, we visualize the proportional relationship of the source dataset in Figure 6 (a). Furthermore, the question types in our training dataset in visualized in Figure 6 (b), and the selected data types are relatively balanced, which enhances the generalizability and robustness of the ranking network in AutoV.
We then generate various visual prompts based on the above source data. Here, we simply use the attention prompting method employed by the API [55] for high-quality generation. To streamline the process, we empirically generate heat maps of attention maps corresponding to CLIP [40] layers , , , and under its setting on the image as our training visual prompts. A specific example (including JSON dictionary and images) can be found in Figure 7. As a result, one question will correspond to four visual prompts, with a total of AutoV being trained on K visual prompts.
Appendix C Evaluation Benchmarks
We conducted experiments on several widely used visual understanding benchmarks.
MMMU. [58] The MMMU is designed to evaluate large vision-language models across a wide range of academic disciplines and complex tasks. It covers over 30 subjects including history, medicine, physics, and more, requiring multimodal comprehension, fine-grained reasoning, and domain-specific knowledge. Each question in MMMU is paired with an image and a multiple-choice question, testing the model’s ability to integrate visual understanding with deep textual reasoning.
GQA. [19] The GQA benchmark is composed of three parts: scene graphs, questions, and images. The image part contains images, as well as the spatial features of images and the features of all objects in images. The questions in GQA are designed to test the understanding of visual scenes and the ability to reason about different aspects of an image.
POPE. [27] The POPE benchmark is primarily used to evaluate the degree of Object Hallucination in models. It reformulates hallucination evaluation by requiring the model to answer a series of specific binary questions regarding the presence of objects in images. Accuracy, Recall, Precision, and F1 Score are effectively employed as reliable evaluation metrics to precisely measure the model’s hallucination level under three different sampling strategies.
TextVQA. [44] The TextVQA benchmark focuses on the comprehensive integration of diverse text information within images. It meticulously evaluates the model’s text understanding and reasoning abilities through a series of visual question-answering tasks with rich textual information. Models need to not only understand the visual content of the images but also be able to read and reason about the text within the images to answer the questions accurately.
MMVet. [57] The MMVet benchmark is designed based on the insight that the intriguing ability to solve complicated tasks is often achieved by a generalist model being able to integrate different core vision-language capabilities. MM-Vet defines 6 core VL capabilities and examines the 16 integrations of interest derived from the capability combination.
LLaVA-In-The-Wild. [31] The LLaVA-In-The-Wild benchmark is built to assess the generalization ability of vision-language models in real-world, unconstrained settings. Unlike curated benchmarks, it features user-collected image-question pairs from diverse sources and domains. This setup presents models with noisy, complex, and unpredictable visual content, offering a realistic measure of their robustness, grounding capability, and performance in open-ended scenarios.
Appendix D Implementation Details.
(1) Data generation stage: LLaVA-v1.5 is consistently adopted as the reference LVLM for generating loss due to resource limitations. (2) AutoV training stage: LLaVA series are trained using DeepSpeed [42] ZeRO2 with A100-40G GPUs within epochs. The batch size (with accumulation) is set to and the learning rate is -. For convenience, Qwen2.5-VL and InternVL2 directly adopt the retrieval strategy in AutoV pre-trained from LLaVA-OneVision.
Appendix E Limitations
The introduced AutoV offers a new perspective on retrieving various visual prompts for LVLMs when faced with different queries. However, due to resource constraints, we only generate the training data via LLaVA-7B model [31], and AutoV might yield superior results with more powerful large vision-language models. Besides, to maintain parameter efficiency, we design our approach as a plug-in module for existing LVLMs, requiring no fine-tuning of the model. We note, however, that naively reusing the LLM frozen parameters may lead to suboptimal performance. Finally, due to current limitations in visual prompt quality, our selection is constrained to attention visual prompts [55], resulting in bounded accuracy gains. We believe that when future advances in visual prompt generation, AutoV can further boost their integrated performance.
Appendix F Broader Impacts
Overall, this research has broader impacts on the visual prompt, feature retrieval, and future development of LVLMs, fostering progress and advancements in the field of vision-language models.
Advancing Method: We propose the AutoV model that learns to automatically select the optimal visual prompt from various candidates based on given textual queries and the input image.
Novel Insights: We are the first to apply the reward modeling loss from RLHF [41] to optimal visual prompt retrieval and validate its effectiveness compared with the Mixture-of-Experts framework.
Inspiring Future Research: This work advances visual prompt retrieval in LVLMs, establishing a foundation for future research. By framing the prompt selection challenge systematically, we enable subsequent exploration of (1) higher-quality prompt generation techniques and (2) more effective retrieval methodologies, collectively pushing the boundaries of multimodal understanding.
Appendix G More Visualization
G.1 Cases of Training Data
Visualization of a case in our training data is in Figure 7, including a JSON dictionary and images.
G.2 Retrieval Strategy of AutoV
Visualization of the retrieval strategy employed by AutoV is in Figure 5, with a total of ten cases.