### v2

# Loss-Oriented Ranking for Automated Visual Prompting in LVLMs

Abstract
Inspired by text prompts in large language models (LLMs), visual prompts have been explored to enhance the reasoning capabilities of large vision-language models (LVLMs). Current methods design heuristic visual prompts, such as overlaying a text-query-guided attention heatmap on the original input image. However, designing effective prompts manually is challenging and time-consuming, and it often fails to explore the benefits of different visual prompts, leading to sub-optimal performance. To this end, we propose AutoV that learns to automatically select the optimal visual prompt from various candidates based on given textual queries and the input image. To train AutoV, we develop an automatic data collection and labeling pipeline that evaluates various visual prompts with a pre-trained LVLM. We input a set of visual prompts into the LVLM and rank them according to the prediction losses generated by the model. Using the ranking as a supervision signal, we train AutoV to automatically choose the optimal visual prompt from various visual prompts for LVLMs. Experiments indicate that AutoV enhances the performance of various LVLMs across multiple image understanding tasks. For instance, LLaVA-OV with AutoV achieves accuracy gain on VizWiz, and AutoV boosts Qwen2.5-VL by on MMMU, highlighting its potential as an optimal visual prompting method.
1 Introduction
Recent advancements in large language models (LLMs) [brown2020language, achiam2023gpt, touvron2023llama, peng2023instruction, bi2024deepseek] have spurred significant progress in large vision-language models (LVLMs), which connect visual encoders with language model decoders to effectively adapt textual reasoning capabilities to visual understanding tasks [liu2023improvedllava, Qwen-VL, chen2023internvl, bai2025qwen2, li2024llava]. Therefore, visual input plays a crucial role in LVLMs, enabling precise perception and fine-grained grounding across various multimodal scenarios.
To further improve the perception and reasoning of LVLMs, visual prompting [api, lin2024training, yang2023fine, shtedritski2023does] has emerged as an important technique analogous to textual prompting in LLMs [kojima2022large, zhou2022cocoop, wei2022chain]. Visual prompts enhance model performance by adding visual annotations such as masks, colorful circles, or heatmaps, thus guiding the attention of LVLMs toward crucial image regions of interest. However, existing visual prompting methods predominantly rely on heuristically designed prompts, limiting their ability to fully exploit the inherent diversity of visual prompts. As illustrated in Figure 1, we list the benchmark-level performance of existing heuristic prompt methods (e.g., FGVP [yang2023fine], Red Circle [shtedritski2023does], and API [api]). While all three methods demonstrate effectiveness on logic-intensive MMMU tasks, both blur-based FGVP and bounding prompts adversely affect vision-centric benchmarks (e.g., VizWiz [bigham2010vizwiz]) and OCR-dependent tasks (e.g., TextVQA [singh2019towards]). This is because heuristic prompt methods fail to automatically tune hyperparameters for each benchmark or even each instance.
Hence, to develop a more effective visual prompt strategy suitable for diverse vision-language tasks, it is vital to adaptively retrieve the optimal visual prompt that most significantly improves the perception and reasoning of LVLMs in response to specific queries. Motivated by this, we introduce AutoV, an innovative visual prompting retrieval framework designed to dynamically select the optimal visual prompt from a set of candidates, explicitly tailored to the textual query and input image. Specifically, we first generate various visual prompt candidates and encode their representation. This diversity enables the framework to capture distinct visual representations that may be optimal for various query-image pairs. Next, we design a lightweight ranking network that efficiently integrates visual and textual information to predict the rankings preferences over prompt candidates. Finally, the highest-ranking candidate is chosen as the input for LVLMs during inference.
To train the ranking network, we establish an automated pipeline for collecting and labeling training data, evaluating the rank of each visual prompt candidate based on the prediction losses of a pre-trained LVLM. Subsequently, the ranking network is supervised using a reward loss that differentiates between selected and rejected candidates. This strategy provides reliable supervision, allowing AutoV to attain adaptive, query-specific prompt retrieval without the need for extensive manual annotation work.
AutoV achieves notable performance improvements across various prominent vision-language models and diverse image understanding benchmarks, including fine-grained comprehension, real-world visual assistance, and complex open-ended reasoning tasks. For instance, integrating AutoV with LLaVA-OneVision [li2024llava] and InternVL2 [chen2023internvl] leads to improvements of and on VizWiz [bigham2010vizwiz], respectively. Additionally, AutoV enhances Qwen2.5-VL [Qwen-VL] on MMMU [yue2024mmmu] by , highlighting its significant advantage in visual prompting reasoning retrieval. Furthermore, AutoV generalizes well, maintaining stable performance across captioning, grounding, and classification tasks. In summary, our contributions are as follows:
We introduce AutoV, a novel automatic visual prompting framework for large vision-language models that adaptively retrieves the optimal visual prompt from a diverse candidate pool, explicitly tailored to each textual query.
We create a scalable and totally automated data collection pipeline equipped with a reward-based loss to effectively train a lightweight ranking network, facilitating accurate and adaptive visual prompt selection without the need for extensive manual annotations.
We perform comprehensive experiments to validate the efficacy and robustness of AutoV across various multimodal models and tasks, showcasing consistent and significant improvements over existing visual prompting methods. Once trained, AutoV integrates seamlessly into various LVLMs without additional fine-tuning.
2 Related Work
Large Vision-Language Models. Recent advances in large language models [radford2019language, brown2020language, achiam2023gpt, touvron2023llama, peng2023instruction, bi2024deepseek] have catalyzed a new wave of large vision-language models [liu2023improvedllava, alayrac2022flamingo, Qwen-VL, chen2023internvl, li2024mini, li2023blip]. By coupling a high-capacity visual encoder with an LLM decoder, LVLMs transfer textual reasoning to visual understanding, and the quality of input images determines the lower bound of multimodal abilities. For instance, the popular LLaVA-OneVision [li2024llava] and Qwen2.5-VL [bai2025qwen2] adopt a higher and dynamic resolution recipe for training to enhance their fine-grained reasoning, while vision-centric scaling in InternVL2 [cai2024internlm2] enlarges the ViT [dosovitskiy2020image] backbone to boost grounding capability. However, the above methods require training times of several hundred GPU days. Here, can we further boost the existing LVLMs by optimizing the visual input itself during the inference stage? The answer is Yes. Visual prompting [sun2023alphaclip, lin2024drawandunderstand, jiang2024joint, zhang2024exploring, api, yang2023fine, shtedritski2023does] is an emerging and effective technique in this direction.
Visual Prompting for LVLMs. LVLMs relying solely on textual inputs exhibited two key limitations: visual hallucinations [bai2024hallucination, huang2024visual] and language-driven biases [wu2024commit, qu2024unified]. To better mitigate these challenges, visual prompting [zhang2023vpgtrans, sun2023alphaclip, lin2024drawandunderstand, wu2024visual] has emerged as an alternative way of supplying cues in the visual modality. Compared with textual prompts, it is more concise and explicit, while allowing guidance that reaches fine-grained, pixel-level details. Early work [shtedritski2023does] shows that simply drawing a red circle around a target object steers CLIP [CLIP] attention, yielding great performance on referring-expression and keypoint-localization benchmarks. FGVP [yang2023fine] extends this idea with fine-grained prompting, replacing coarse shapes with blur-reversal prompts based on segmentation masks, further improving accuracy on RefCOCO [zhang2019referring]. AlphaCLIP [sun2023alphaclip] trains an augmented version of CLIP with an auxiliary alpha channel for indicating regions at the expense of increased training overhead. API [api] proposes attention prompting on images, overlaying text-guided attention heatmaps, boosting LVLMs on VQA, and open-ended reasoning tasks. Although previous studies have explored various visual prompts, it remains unclear which prompt is optimal for a specific image-question pair, leading to a decrease in the practicability of visual prompts. To address this limitation, we introduce our visual prompting method that learns to retrieve the optimal visual prompt for a specific image-question pair.
Visual Retrieval for LVLMs. A variety of strategies have been proposed for visual retrievals, such as visual token selection [chen2024image, zhang2024sparsevlm, yang2024visionzip] for LVLMs. Existing visual prompt selection pipelines generally adhere to three main paradigms. (1) The heuristic paradigm identifies benchmark-level optimal parameters through extensive evaluation without instance-specific adaptation. (2) The random paradigm offers a baseline characterized by instance-level variation. (3) The Mixture-of-Experts (MoE) paradigm [fedus2022switch, bao2022vlmo, zong2024mova] provides the most direct form of adaptation by utilizing the GateNet specifically for visual retrieval, which is susceptible to overfitting during training.
Notably, our AutoV fundamentally differs from retrieval-augmented methods [qi2024rora, xing2025re]. Unlike them, which retrieve supplementary visual data from external databases and require model finetuning, AutoV selects the most suitable visual prompt based solely on the internal features of the current image through a lightweight ranking network. Moreover, it operates within a compact prompt pool rather than a large-scale database, yielding substantial efficiency gains.
3 Proposed Approach: AutoV
In this section, we propose the AutoV for optimal visual prompt retrieval. Our framework includes four critical components: (1) feature extraction of prompt candidates, (2) candidate ranking consisting of modality interaction and single-modality mapping modules, (3) reward loss supervision along with automated data generation pipeline, and (4) an efficient and robust inference pipeline. The overall architecture is illustrated in Figure 2, where the four modules are color-coded for clarity. Best viewed in color.
3.1 Feature Extraction of Prompt Candidates
The primary goal is to effectively choose the most appropriate visual prompt in response to various text queries and input images. For a group of input visual prompts , we employ a visual encoder (e.g., like CLIP [CLIP]) to generate the visual feature:
Next, we apply a projection matrix to transform each visual feature into language embedding tokens , aligning the same dimensionality as the word embedding space in the language model. As a result, we obtain a set of visual candidates represented in the form of visual tokens to rank, where denotes the length of vision tokens and is the language embedding dimension. Notably, both the visual encoder and the alignment projector are directly inherited from a fully trained LVLM, and we adopt them without any additional fine-tuning.
3.2 Candidate Ranking Network
The goal of the ranking network is to effectively rank the visual prompt candidates based on their relevance to the textual query, by modeling the relationship between the visual tokens and the input text. Therefore, we first propose the modality interaction module to enable visual tokens to fuse the context information of text tokens. Inspired by observations in [zhang2025llava] that modality-fusion occurs in the early layers of the LLM, we use the first layer of the LLM decoder for modality interaction. Thus, visual candidate tokens and text tokens are concatenated and fed into the interaction module:
with the outputs are and . Next, we input visual candidate tokens into a vision mapping module, which is implemented using a feed-forward network (FFN). The generated final candidate features are , where is the output dimension of the vision mapping module to save computational cost. Besides, the query tokens are also converted by a text mapping module to get the query embeddings . In AutoV, we define visual retrieval as a customized ranking task rather than a classification or regression problem, and the reason is discussed in Section 3.3. Therefore, we calculate the reward scalar for given visual candidate tokens , where the context similarity with the query serves as the quantitative target. We utilize a cross-attention operation to achieve:
where denotes the th visual prompt candidate. Here, only the vision and text mapping modules need additional training, each consisting of a single FFN with two linear layers. The modules comprise parameters in total, thereby resulting in lower training costs.
3.3 Reward Loss Supervision
Automated Data Generation. We create customized data to train the ranking network to select the optimal visual prompt for a given LVLM. Our approach is grounded in a simple intuition: a superior visual prompt should yield lower language modeling loss when input into any trained LVLM. We optimize the ranking network using relative loss instead of a one-hot selection vector, as correctness can stem from the LVLM’s language priors without relying on any input content [xiao2024can, yu2024frame]. In contrast, the inference loss indicates the confidence of models in generating answers based on the input visual-question pair: a lower loss indicates higher confidence [guo2017calibration, zhang2024unveiling].
For each group of visual candidates, we generate (VP, query) pairs and comprehensively evaluate them. Each pair is processed by a fully trained vision-language model to compute its combination loss, defined as the modeling loss relative to the ground truth, and sorted to get a rank. To ensure high-quality ranking data, we implement two filtering criteria for (VP, query) pairs: (1) we remove pairs exhibiting low loss variance, as their insensitivity to visual prompts suggests answers are generated from the language priors rather than actual input content; (2) we exclude pairs with excessively high average loss, which indicates outlier case or weak visual prompt-query correlation.
For each raw (image, query) pair, we generate a training sample structured as a quadruple:
with an illustrative example provided in Figure 2, and the dataset we adopt is detailed in Appendix A.
Reward Loss Supervision. Inspired by the reward modeling from human feedback in OpenAI [ouyang2022training], we train our ranking network using reward modeling to align its partial order with the optimal visual prompt. Here, we model visual retrieval as a customized ranking task rather than a classification or regression problem. This is because the objective is not to assign absolute labels or scores, but to identify the most appropriate prompt from a set of candidates by comparing their relative effectiveness in enhancing multimodal understanding. Ranking allows the model to learn fine-grained preferences across diverse query-image combinations and better align with the real-world setting.
In Figure 3, we pair all visual prompt candidates exhaustively, ensuring each prompt is combined with every other prompt exactly once. This results in generating training pairs for each instance. The reason is that, unlike pointwise scoring, which assigns low scores to all responses for inherently challenging tasks, pairwise comparison allows for finer discrimination between responses. Based on the rank and loss from the quadruples, we reinforce the visual prompt with lower loss (chosen sample) for each pair, while penalizing the higher-loss visual prompt (rejected sample).
Specifically, the reward loss is designed as:
where is the scalar output of the ranking network for the chosen visual prompt out of the pair, represents the rejected one, and denotes the Sigmoid function.
3.4 Robust Inference Pipeline
During the inference, we input multiple candidate visual prompts along with the query into the large vision-language model and pass them through our ranking network, as shown in Figure 2. The network outputs a reward scalar for each visual prompt, and we select the highest-scoring one. Its visual tokens are concatenated with the text embeddings and decoded by the LLM to generate the final answer.
Besides, to reduce distributional bias, we introduce a pre-filtering step: the prompt whose visual features are farthest from other candidates, as measured by cosine distance, is removed in advance, further mitigating the negative effects of suboptimal prompts and enhancing robustness.
4 Experiments
4.1 Experiment Settings
Models. We apply our AutoV to four mainstream open-source large vision-language models, i.e., LLaVA-1.5 series [liu2024visual], LLaVA-OneVision [li2024llava], Qwen2.5-VL [bai2025qwen2], and InternVL2 [chen2024expanding]. In particular, LLaVA-1.5 series build on CLIP-ViT [CLIP] and Llama2 [touvron2023llama2]. LLaVA-OneVision enhances it with SigLIP [siglip] and Qwen2 [bai2023qwen]. Qwen2.5-VL introduces a dynamic-resolution encoder, and InternVL2 integrates InternViT [chen2023internvl] with InternLM2 [cai2024internlm2].
Benchmarks. AutoV is extensively evaluated on fifteen vision-language benchmarks, covering video reasoning (CVBench [tong2024cambrian]), vision-centric tests (BLINK [fu2024blink], MMVP [tong2024eyes], MMStar [chen2024we]), comprehensive multimodal evaluation (MM-Vet [yu2023mm]), knowledge/math reasoning (MMMU [yue2024mmmu], MathVista [lu2023mathvista]), robustness and VQA (VizWiz [bigham2010vizwiz], Real-World QA [RealworldQA], TextVQA [singh2019towards], LLaVA-Wild [liu2023improvedllava]), grounding (RefCOCO+ [yu2016modeling]), object-hallucination probing (POPE [li2023evaluating]), captioning (COCO-Caption2017 [lin2014microsoft]), and multimodal classification (ImageNet-1K [deng2009imagenet]).
4.2 Main Results
| Method | MMMU | POPE | VizWiz | MMVet | VQA | LLaVA | MMStar | MathVista | Avg. |
| LLaVA-v1.5 (7B) | |||||||||
| Base Model | 36.3 | 85.9 | 50.0 | 30.6 | 58.2 | 65.4 | 33.2 | 34.2 | |
| FGVP (NIPS23) | 36.9 | 72.6 | 47.5 | 29.0 | 46.3 | 55.6 | 33.0 | 34.6 | |
| Circle (ICCV23) | 37.0 | 84.5 | 49.6 | 31.3 | 55.5 | 62.2 | 33.4 | 35.0 | |
| API (ECCV24) | 37.4 | 86.6 | 51.3 | 31.8 | 58.4 | 67.2 | 34.0 | 35.7 | |
| AutoV | 38.7 | 87.4 | 52.1 | 32.9 | 60.6 | 68.6 | 35.2 | 36.9 | |
| LLaVA-v1.5 (13B) | |||||||||
| Base Model | 36.7 | 85.9 | 53.6 | 36.1 | 61.2 | 66.6 | 32.8 | 35.0 | |
| FGVP (NIPS23) | 36.9 | 76.2 | 46.8 | 29.7 | 49.2 | 63.7 | 31.9 | 35.3 | |
| Circle (ICCV23) | 36.9 | 84.6 | 53.8 | 31.2 | 58.1 | 66.9 | 33.5 | 36.2 | |
| API (ECCV24) | 37.3 | 86.8 | 54.6 | 37.6 | 62.1 | 67.6 | 33.0 | 37.0 | |
| AutoV | 38.9 | 87.8 | 56.3 | 38.0 | 62.7 | 69.5 | 34.2 | 38.3 | |
| LLaVA-OneVision (7B) | |||||||||
| Base Model | 49.4 | 89.1 | 58.2 | 48.8 | 76.1 | 80.6 | 61.8 | 63.2 | |
| FGVP (NIPS23) | 49.6 | 82.9 | 58.9 | 49.1 | 65.4 | 70.8 | 58.8 | 63.4 | |
| Circle (ICCV23) | 50.4 | 88.2 | 60.2 | 47.2 | 77.2 | 80.8 | 62.7 | 65.4 | |
| API (ECCV24) | 50.8 | 90.3 | 66.9 | 51.5 | 77.7 | 81.9 | 63.0 | 64.9 | |
| AutoV | 54.0 | 91.0 | 68.4 | 52.9 | 78.0 | 85.3 | 64.5 | 66.8 | |
| InternVL2 (8B) | |||||||||
| Base Model | 48.6 | 87.8 | 58.6 | 46.5 | 77.0 | 74.6 | 62.0 | 58.3 | |
| FGVP (NIPS23) | 49.6 | 84.2 | 58.3 | 44.3 | 70.7 | 75.7 | 60.0 | 58.9 | |
| Circle (ICCV23) | 48.3 | 86.0 | 57.5 | 45.8 | 74.9 | 75.4 | 63.4 | 58.4 | |
| API (ECCV24) | 50.5 | 88.9 | 61.2 | 48.6 | 77.0 | 76.2 | 63.9 | 59.4 | |
| AutoV | 51.7 | 89.6 | 64.8 | 50.2 | 78.0 | 79.5 | 65.8 | 70.5 | |
| Qwen2.5-VL (7B) | |||||||||
| Base Model | 50.1 | 87.5 | 69.5 | 56.6 | 83.0 | 98.0 | 62.4 | 68.2 | |
| FGVP (NIPS23) | 49.6 | 81.6 | 70.8 | 43.1 | 78.0 | 98.2 | 63.4 | 68.0 | |
| Circle (ICCV23) | 51.6 | 84.2 | 69.2 | 44.5 | 80.5 | 97.5 | 65.1 | 69.1 | |
| API (ECCV24) | 51.0 | 87.9 | 72.1 | 58.0 | 85.3 | 100.6 | 63.0 | 67.6 | |
| AutoV | 53.9 | 89.5 | 74.4 | 59.1 | 86.1 | 102.3 | 65.6 | 70.2 | |
Our results are listed in Table 1, and we compare AutoV with three previous visual prompting methods designed for LVLMs, including FGVP [yang2023fine], RedCircle [shtedritski2023does], and API [api]. The training data and implementation details can be found in the Appendix A and C, respectively. Here, we employ the above three types of visual prompts111The pool includes four attention-based visual prompts extracted from different transformer layers of CLIP in API [api], together with two additional prompts of distinct types, resulting in six images in total. as the candidate pool for AutoV to retrieve from optimally.
The four main observations from the experimental results are as follows: (1) Our AutoV significantly outperforms existing methods, demonstrating consistent advantages across all models and benchmarks. Specifically, for LLaVA-7B, LLaVA-13B, LLaVA-OneVision, InternVL2, and Qwen2.5-VL, the average improvements relative to the base model are , , , and , respectively. Our method performs particularly well for LLaVA-OneVision on VizWiz, with an improvement of . Besides, our method appears to be more effective and robust on the VQA task, where other methods often yield negative gains. (2) Our approach shows the most significant performance advantage on MMMU compared to the API. With an average score of for our method versus for theirs, the relative difference of underscores the importance of accurate case-level visual cues retrieval instead of benchmark-level visual prompts selection. The other two visual prompting methods, lacking the ability to adapt to different queries, do not perform well on reasoning tasks. (3) Our method further supplements inherently less accurate models. For instance, LLaVA-7B equipped with our approach achieves higher accuracy than the 13B-scale model on 5 out of 8 tasks. This means that AutoV can be utilized for model efficiency on certain specific tasks: a few additional MLP layers are a better choice compared with hundreds of times larger parameter sizes. (4) Both Qwen2.5-VL and InternVL2 integrate the retrieval optimal prompting strategy from AutoV, which is pre-trained on LLaVA-OneVision. This integration results in an average accuracy gain of . Notably, the improvements are consistent and observed in all individual tasks, including complex reasoning (e.g., MathVista) and vision-centric scenarios (e.g., MMStar). This demonstrates that the effectiveness of the retrieval-enhanced prompting strategy is model-agnostic and generalizes well across heterogeneous LVLM architectures, regardless of pretraining scales or backbone variations.
Additionally, we extend the evaluation to a broader range of vision-language benchmarks, including CVBench, BLINK, Real-World QA, and MMVP, as well as classic tasks such as image captioning, visual grounding, and multimodal classification. We conduct experiments using both LLaVA-OneVision and Qwen2.5-VL as base models to examine the generalization capability of AutoV. The detailed quantitative results are provided in the Appendix D.
4.3 Analysis
4.3.1 Comparison with Other Retrieval Methods
| Method | MMMU | VizWiz | MMVet | Avg. |
| Baseline | 36.3 | 50.0 | 30.6 | 39.0 |
| only | 36.7 | 50.7 | 31.8 | 39.8 |
| only | 36.9 | 51.3 | 31.4 | 39.9 |
| only | 37.4 | 50.6 | 31.7 | 39.9 |
| only | 37.2 | 51.2 | 31.2 | 39.9 |
| Random | 37.1 | 50.7 | 31.3 | 39.7 |
| MoE | 37.6 | 50.7 | 32.0 | 40.1 |
| AutoV | 38.7 | 52.1 | 32.9 | 41.3 |
To analyze why ranking is more suitable for retrieving visual prompts, we compare it with three other retrieval methods. The experiments are conducted on LLaVA-7B. The retrieval pool is identical to that used in the main experiments.
Heuristic Fixed Retrieval. It involves a single fixed type of visual prompts, based on its predefined rules or heuristics. As shown in rows 3 to 6 of Table 2, we find that each visual prompt could potentially be the optimal solution for different benchmarks. For instance, achieves the best results on MMMU, while is optimal on MMVet, and there is a significant variation in the performance of different types of visual prompts across various benchmarks. In summary, heuristic retrieval offers simplicity in retrieval processes with the need for conducting benchmark-level multiple rounds of trials, consuming more resources.
Random Retrieval. It picks up one visual prompt from the candidate pool without any specific criteria or order. The th row of Table 2 lists its results. By comparing with the individual visual prompt, the performance of random retrieval is on par with or slightly superior across the three benchmarks. For instance, the random algorithm outperforms in MMMU by , in MMVet by . It demonstrates that by flexibly selecting from multiple visual prompt candidates, even random retrieval has the potential to outperform individual visual prompts, and the conclusion provides valuable insights for our approach.
MoE (or GateNet) Retrieval. It is a common technique for selecting the most relevant expert during model decision-making, where Gumbel Softmax enables differentiable weighting of the gating network [fedus2022switch]. In our scenario, the features of multiple visual prompts are input to GateNet to determine which one should be retained and fed into the LLM. We insert GateNet into the output of the projection layer and train it with the same amount of data and training time as AutoV. The results can be seen in the th row of Table 2. We observe a improvement with MoE compared to the baseline. However, the improvement compared to the random algorithm was only , lagging significantly behind AutoV by . We suspect that the Mixture-of-Experts architecture encounters a lack of generalization ability during the fine-tuning stage, which can easily lead to overfitting. Therefore, MoE retrieval may not be the optimal approach for visual prompt retrieval in LVLMs. It is important to adopt a more robust framework that integrates textual guidance to enhance overall performance effectively.
Retrieval Visualization. We visualize the prompt distribution of retrieval strategies within - for significance. Figure 4 presents a bar graph comparing the retrieval decisions of GateNet and our AutoV on MMVet. The bar heights indicate the frequency of visual prompt retrieval, while the overlaid line plots depict the independent accuracy for each prompt. Our analysis reveals a notable discrepancy between the GateNet retrieval and the actual inference accuracy of visual prompts. For instance, GateNet retrieves , the second-most useful prompt, only times (38/218, of cases), despite its high accuracy. In contrast, AutoV exhibits stronger alignment with the standalone accuracy of visual prompts, and prefers to select and , the two prompts that own superior accuracy.
4.3.2 Scaling Law in the Visual Prompts Pool
| Pool Size | 0 | 1 | 2 | 3 | 4 | 5 | 6 | 7 | 8 |
| Source | Baseline | +{API-L} | +{API-L} | +{API-L} | +{API-L} | +{RedCircle} | +{FGVP} | +{VTPrompt} | +{TVP} |
| MMMU | 36.3 | 37.3 | 37.5 | 38.3 | 38.3 | 38.6 | 38.7 | 38.7 | 38.9 |
| VizWiz | 50.0 | 50.3 | 50.8 | 51.0 | 51.9 | 52.1 | 52.1 | 52.6 | 52.6 |
| MMVet | 30.6 | 31.8 | 32.0 | 32.6 | 32.8 | 32.9 | 32.9 | 33.1 | 33.1 |
| Avg. | 39.0 | 39.8 | 40.1 | 40.7 | 41.0 | 41.2 | 41.3 | 41.5 | 41.6 |
Having more visual prompt candidates can offer a more diverse selection, enhancing the flexibility and adaptability of LVLMs. However, is it always better to have more visual prompts in AutoV? In the main experiment, we used the attention prompting images [api] from layers , , , and in the CLIP [CLIP] model, FGVP [yang2023fine], and RedCircle [shtedritski2023does] as inputs, with a total of visual prompts. Here, we consider extra scenarios ranging from prompts and conduct experiments on the MMMU, VizWiz, and MMVet. The results are listed in Table 3. First, we observe that when the visual prompts themselves can lead to performance improvement, the accuracy increases as their quantity grows, but the rate of accuracy improvement slows down as the number continues to rise. Taking the experiments on MMMU as an example, as the number of prompts increases, the performance gain remains at when increasing from to prompts. However, when further increasing to prompts, the improvement drops to . Then, even if a single visual prompt itself degrades accuracy, incorporating it into the candidate pool does not hurt overall performance and can even lead to slight improvements, thanks to the inherent retrieval capability of AutoV. Taking the results on VizWiz as an example: although both RedCircle and FGVP individually show negative gains ( and , respectively), including them in the candidate pool leads to an overall improvement. In summary, AutoV exhibits notable robustness to variations in visual prompt quality, consistently preserving or enhancing overall performance across diverse settings. We recommend selecting known high-quality visual prompts as candidates, while limiting to fewer than instances to strike the balance between performance and efficiency.
4.3.3 The Transferability of AutoV
| Models | VizWiz | LLaVA |
| Gemini-1.5-Pro | 45.9 | 77.9 |
| +AutoV | 55.5 | 81.1 |
| GPT-4o | 52.8 | 68.0 |
| +AutoV | 61.8 | 70.9 |
| Qwen-VL-Max | 64.6 | 85.8 |
| +AutoV | 71.9 | 88.4 |
The retrieval strategy in our AutoV, trained on one LVLM, can be transferred to other LVLMs. First, our method demonstrates broad applicability across open-source models. In Table 1, both InternVL2 and Qwen2.5-VL achieve a improvement by directly adopting the retrieval strategy from AutoV pre-trained on LLaVA-OneVision. Besides, Table 4 reveals that even closed-source models like Gemini-Pro [team2024gemini] and GPT-4o [hurst2024gpt] show substantial gains, averaging and points respectively. Therefore, AutoV owns good transferability, as its retrieval strategy consistently boosts performance across both open- and closed-source LVLMs without model-specific adaptation.
4.3.4 Statistical Significance Analysis
To rigorously evaluate the reliability of our reported improvements, we conduct significance testing on the results in Table 1 (LLaVA-1.5 7B) across multiple random seeds. Taking the MMMU benchmark as an example, a paired t-test [hsu2014paired] confirms that improvement of AutoV over the baseline is highly significant (mean , ), with an extremely small -value, indicating that the observed gain is unlikely to arise from random chance. This supports our claim that AutoV consistently enhances VLM performance across repeated runs. Detailed statistical tables and analyses are provided in the Appendix E.
4.3.5 The Efficiency of AutoV
While we enhance LVLM performance through AutoV, we uphold efficiency optimizations to ensure practicality. Initially, training a high-quality AutoV on the LLaVA-7B model within epochs only requires hours utilizing NVIDIA A-GB GPUs. Furthermore, because of its transferability, training additional AutoV instances for each vision-language model is unnecessary, and we can generate the training samples via smaller models. Lastly, although AutoV introduces additional TFLOPs during inference (e.g., multiple image pre-processing and visual feature extraction), we can reduce latency by downsizing the final similarity matrix and single-prompt selection before LLM processing. As a result, the inference latency of the LLaVA-7B model increases by just ms (from ms to ms), while yielding a gain in average accuracy.
4.3.6 Qualitative Visualization
To validate the effectiveness of our retrieval strategy, we visualize an example sourced from LLaVA-Bench within {API-L-Dark, RedCircle, API-L-White, FGVP}. As shown in Figure 5, we list the output results of AutoV when we enter different queries. When faced with a question that focuses on the global context information, AutoV chooses (API-L-Dark), where the black mask removes the background. While referring to the ”intended effect” in the image, AutoV selects (RedCircle), which directly uses red circles to highlight the weird spots. More visualization retrieval cases can be found in the Appendix F.
5 Conclusion
In this paper, we introduce AutoV, an adaptive visual prompting framework that retrieves optimal visual prompts tailored to textual queries and image contexts for large vision-language models. By leveraging a lightweight ranking network trained with reward-based supervision, AutoV consistently outperforms existing heuristic prompting methods. Extensive experiments across multiple LVLM architectures and benchmarks demonstrate its effectiveness and generality. Overall, AutoV provides a practical and robust solution for enhancing LVLMs’ performance through adaptive visual prompting. Future work explores integrating AutoV with more advanced prompt generation techniques or extending its framework to broader multimodal settings beyond vision-language tasks.
Supplementary Material
Contents
A Details of Training Data
| Method | CVBench | BLINK | RealWorldQA | MMVP | COCOCap | RefCOCO+ | ImageNet-1K | Avg. |
| LLaVA-OneVision (7B) | ||||||||
| Base Model | 79.6 | 48.2 | 66.0 | 60.7 | 61.1 | 68.8 | 92.3 | |
| API (ECCV24) | 80.7 | 49.4 | 66.4 | 61.5 | 61.9 | 69.4 | 92.6 | |
| AutoV | 83.0 | 51.6 | 68.9 | 64.2 | 63.7 | 71.8 | 95.0 | |
| Qwen2.5-VL (7B) | ||||||||
| Base Model | 80.6 | 56.4 | 68.5 | 61.3 | 40.7 | 72.6 | 94.7 | |
| API (ECCV24) | 81.1 | 57.3 | 69.2 | 61.7 | 41.4 | 73.5 | 95.4 | |
| AutoV | 83.0 | 59.0 | 71.1 | 64.3 | 43.3 | 75.2 | 97.1 | |
A.1 Attention Visual Prompts for Training
As vision foundation models advance rapidly [CLIP, oquab2023dinov2, siglip, kirillov2023segment], pretrained on language-image pairs, they serve as the visual encoders for large vision-language models. A representative model is CLIP [CLIP], , which consists of a vision encoder and a text encoder, mapping the image and text prompt into the aligned latent space. Therefore, we can directly obtain the similarity between vision modality and text modality feature with .
Inspired by the idea that text-image attention (or similarity ) can be visualized as visual prompts for LVLMs [api, lin2024training], where similarity heatmaps overlaid on the image highlight key regions requiring greater attention for prompt-based reasoning. In this work, we adopt a fine-grained similarity visualization manner in API [api], with the -th layer similarity defined as follows:
where denotes the number of transformer layers, is the projection layer with a fully-connected layer and normalization operation, and MSA stands for multi-head self-attention structure. is input vision tokens for the -th layer, while indicates the class token in sequence .
A.2 The Composition of Training Data
Our training source dataset is composed of several components used in LLaVA [liu2023improvedllava] to ensure that AutoV is trained on a diverse range of image-question pairs. It includes K samples from the COCO train dataset [lin2014microsoft], K samples from the Visual Genome (VG) dataset [krishna2017visual], K samples from the GQA dataset [hudson2019gqa], K samples from the OCRVQA dataset [mishra2019ocr], and K samples from the TextVQA dataset [singh2019towards]. In total, the source data contains K image-question pairs. Here, we visualize the proportional relationship of the source dataset in Figure 6 (a). Furthermore, the question types in our training dataset in visualized in Figure 6 (b), and the selected data types are relatively balanced, which enhances the generalizability and robustness of the ranking network in AutoV.
We then generate various visual prompts based on the above source data. Here, we simply use the attention prompting method employed by the API [api] for high-quality generation. To streamline the process, we empirically generate heat maps of attention maps corresponding to CLIP [CLIP] layers , , , and under its setting on the image as our training visual prompts.
A specific example (including JSON dictionary and images) can be found in Figure 7. As a result, one question will correspond to four visual prompts, with a total of AutoV being trained on K visual prompts.
B Evaluation Benchmarks
We conduct comprehensive experiments on fifteen widely used vision-language benchmarks.
MMMU. [yue2024mmmu] The MMMU is designed to evaluate LVLMs across a wide range of academic disciplines and complex tasks. It covers over subjects, including history, medicine, physics, and more, requiring multimodal comprehension, fine-grained reasoning, and domain-specific knowledge. Each question is paired with an image and a multiple-choice question, testing the ability to integrate visual understanding with deep textual reasoning.
GQA. [hudson2019gqa] The GQA benchmark is composed of three parts: scene graphs, questions, and images. The image part contains images, as well as the spatial features of images and the features of all objects in images. The questions in GQA are designed to test the understanding of visual scenes and the ability to reason about different aspects of an image.
POPE. [li2023evaluating] The POPE benchmark is primarily used to evaluate the degree of Object Hallucination in models. It reformulates hallucination evaluation by requiring the model to answer a series of specific binary questions regarding the presence of objects in images. Accuracy, Recall, Precision, and F1 Score are effectively employed as reliable evaluation metrics to precisely measure the model’s hallucination level under three different sampling strategies.
TextVQA. [singh2019towards] The TextVQA benchmark focuses on the comprehensive integration of diverse text information within images. It meticulously evaluates the model’s text understanding and reasoning abilities through a series of visual question-answering tasks with rich textual information. Models need to not only understand the visual content of the images but also be able to read and reason about the text within the images to answer the questions accurately.
MMVet. [yu2023mm] The MMVet benchmark is designed based on the insight that the intriguing ability to solve complicated tasks is often achieved by a generalist model that can integrate different core vision-language capabilities. MM-Vet defines 6 core capabilities and examines the 16 integrations of interest derived from the capability combination.
LLaVA-In-The-Wild. [liu2023improvedllava] The LLaVA-In-The-Wild benchmark is to assess the generalization ability of LVLMs in real-world, unconstrained settings. Unlike curated benchmarks, it features user-collected image-question pairs from diverse sources and domains. This setup presents models with noisy, complex, and unpredictable visual content, offering a realistic measure of their robustness, grounding, and performance in open-ended scenarios.
MMStar. [chen2024we] MMStar evaluates six vision-language capabilities across 18 fine-grained dimensions using visually dependent questions. It introduces Multi-modal Gain and Multi-modal Leakage metrics to measure visual contribution and data leakage. This benchmark rigorously tests models’ true visual reasoning and generalization abilities.
MathVista. [lu2023mathvista] MathVista assesses mathematical reasoning grounded in visual contexts such as diagrams and plots. It integrates 28 prior datasets and adds new subsets targeting logical, functional, and scientific figure reasoning. The benchmark challenges models to combine precise visual understanding with compositional calculations and multi-step quantitative reasoning.
CVBench. [cvbench2025] CVBench measures models’ ability to reason across multiple videos. Its tasks involve object association, event association, and complex multi-video reasoning between distinct video clips. The dataset highlights large performance gaps between humans and models in multi-video reasoning and temporal integration.
BLINK. [fu2024blink] BLINK targets core perceptual abilities that humans solve instantly but multimodal models struggle with. It reformulates 14 classic vision tasks into 3,800 multiple-choice questions testing fine-grained perception. Results reveal that current models still lack fundamental visual understanding compared to humans.
Real-World QA. [realworldqa2025] In order to develop useful real-world AI assistants, it is crucial to advance a model’s understanding of the physical world. RealWorldQA is a benchmark just for the design. The dataset consists of anonymized images taken from vehicles, in addition to other real-world images. The initial release of the RealWorldQA consists of over 700 images, with a question and easily verifiable answer for each image.
MMVP. [tong2024eyes] MMVP exposes perceptual blind spots in vision-language models using “CLIP-blind” image pairs. These are pairs of images that contain clear visual differences to humans but are nearly indistinguishable to CLIP-based vision encoders. Leveraging such tricky image pairs, the benchmark pinpoints core weaknesses in visual representation and perception fidelity.
COCOCap. [chen2015microsoftcococaptionsdata] Microsoft COCO Image Captioning benchmark, a standard test of a model’s ability to generate natural language descriptions of images. It contains over 330,000 images with five human-written captions each, scored by metrics like BLEU [papineni-etal-2002-bleu] and CIDEr [vedantam2015ciderconsensusbasedimagedescription]. Performance reflects a model’s capacity for coherent and contextually accurate visual description.
RefCOCO+. [kazemzadeh-etal-2014-referitgame] RefCOCO+ measures models’ ability to ground appearance-based referring expressions in images. Its annotations forbid spatial words, forcing reliance on attributes. This benchmark tests precise visual grounding under complex, attribute-driven conditions.
ImageNet-1K. [imagenet15russakovsky] ImageNet-1K is a foundational benchmark for large-scale image classification. It contains 1,000 object categories with 1.28M training and 50K validation images. Top-1 and Top-5 accuracy serve as key metrics for evaluating visual recognition performance.
C Implementation Details
(1) Data generation stage: LLaVA-v1.5 is consistently adopted as the reference LVLM for generating loss due to resource limitations. (2) AutoV training stage: The LLaVA series is trained using DeepSpeed [rasley2020deepspeed] ZeRO2 with A100-40G GPUs within epochs. The batch size (with accumulation) is set to and the learning rate is -. For convenience, Qwen2.5-VL, InternVL2, Gemini-1.5-Pro, GPT-4o, and Qwen-VL-Max directly adopt the retrieval strategy in AutoV pre-trained from LLaVA-OneVision.
D More Main Experiments
In this section, we present the extended results that complement the main experiments reported in Section 4. Specifically, we provide detailed quantitative results and additional analyses of AutoV on a broader range of benchmarks, including CVBench, BLINK, Real-World QA, and MMVP, as well as classical tasks such as COCO2017 Captioning, RefCOCO+ Grounding, and ImageNet-1K Classification. All experiments are conducted using LLaVA-OneVision 7B and Qwen2.5-VL 7B as base models. The results consistently demonstrate that AutoV effectively enhances multimodal understanding and reasoning across diverse domains, confirming its model-agnostic generalization capability.
| Run | Seed | AutoV | Baseline | Diff. |
| 1 | 10 | 38.7 | 36.3 | +2.4 |
| 2 | 20 | 38.7 | 36.3 | +2.4 |
| 3 | 30 | 38.8 | 36.2 | +2.6 |
| 4 | 40 | 38.7 | 36.2 | +2.5 |
| 5 | 50 | 38.8 | 36.4 | +2.4 |
E Statistical Significance Analysis
To further verify the robustness of the improvement in the main experiments, we conduct significance testing across multiple random seeds on representative benchmarks. Specifically, we present the detailed analysis process for MMMU as an example, while the same manner applies to the remaining benchmarks. The results are shown in Table 6, and the analysis is presented below:
Mean : +2.06%
Std Dev:
-value:
The above analysis provides solid statistical evidence that AutoV delivers consistent and reproducible improvements across runs, highlighting its robustness and reliability in enhancing vision-language model performance.
F More Visualization Results
F.1 Cases of Training Data
Figure 7 shows one training example in our dataset, including its JSON structure and the corresponding attention maps. The JSON entry stores the image path, human–GPT conversation, different visual prompt paths from the API, as well as the predicted ranking and loss values. The right side visualizes these visual prompts, illustrating how different prompts emphasize varying regions of the same image.
F.2 Retrieval Strategy of AutoV
To further demonstrate the effectiveness of our retrieval strategy, Figure 8 provides additional visualizations of ten representative cases beyond those presented in the main text, showing how AutoV retrieves the most suitable visual prompt according to different textual queries. These examples are sourced from different benchmarks, covering multiple visual prompt sets, including {API-L-Dark, RedCircle, API-L-White, FGVP}. Best viewed in color.